OpenStudy (anonymous):

The solution set of 2x^3/4+1=17 is? Fractions like that confuse me ._.

jimthompson5910 (jim_thompson5910):

Is it \[\LARGE 2x^{\frac{3}{4}}+1=17\] or is it \[\LARGE \frac{2x^{3}}{4}+1=17\] ???

OpenStudy (anonymous):

The first one jim

OpenStudy (anonymous):

\[2x^{\frac{3}{4}}+1=17\] \[2x^{\frac{3}{4}}=16\] \[x^{\frac{3}{4}}=8\] is a start

OpenStudy (anonymous):

then \[x=8^{\frac{4}{3}}=\sqrt[3]{8^4}=2^4=16\]

jimthompson5910 (jim_thompson5910):

\[\LARGE 2x^{\frac{3}{4}}+1=17\] \[\LARGE 2\sqrt[4]{x^3}+1=17\] \[\LARGE 2\sqrt[4]{x^3}=16\] \[\LARGE \sqrt[4]{x^3}=8\] \[\LARGE x^3=8^4\] \[\LARGE x^3=(2^3)^4\] \[\LARGE x^3=2^{3*4}\] \[\LARGE x^3=2^{4*3}\] \[\LARGE x^3=(2^4)^3\] \[\LARGE x=\sqrt[3]{(2^4)^3}\] \[\LARGE x=2^4\] \[\LARGE x=16\] So the answer is \[\LARGE x=16\]

OpenStudy (anonymous):

Ahhhh this makes so much sense now. Thank you!